OpenStudy (anonymous):

When n = energy efficiency and Pin = energy input and Pout = energy output, how can you mathematically represent the correct relationship between the energy you put in something and the energy you get out? A. n = P out * P in B. n = P out / P in C. n = P out - P in

OpenStudy (jamesj):

If a machine/process is perfectly energy efficient, then \( \nu = 1 \). That being the case, which is the only definition that works?

OpenStudy (anonymous):

aint that pin/pout=efficiency?

OpenStudy (jamesj):

*correction. Not nu = 1, but eta = 1, \( \eta = 1 \) For the record, if it is absolutely energy inefficient, then \( \eta = 0 \).

OpenStudy (anonymous):

my oh my means i screwed up!!

OpenStudy (anonymous):

yeah but i am having a problem of figuring out the problem

OpenStudy (anonymous):

i mean the answer is fine but can you explain it so that i can figure it out my self next time

OpenStudy (jamesj):

If a process, is perfectly efficient, then Energy out = Energy in. P_out = P_in Therefore \[ \eta = \frac{P_{out}}{P_{in{\}} = 1 \]

OpenStudy (jamesj):

\[ \eta = \frac{P_{out}}{P_{in}} = 1 \]

OpenStudy (anonymous):

oh ok

OpenStudy (jamesj):

If it absolutely inefficient, then P_out = 0 and \[ \eta = \frac{P_{out}}{P_{in}} = 0 \] If half the energy in comes out, then \[ P_{out} = \frac{1}{2} P_{in} \] and therefore \[ \eta = \frac{P_{out}}{P_{in}} = \frac{1}{2} \] So you can see, this is very natural definition of efficiency. The quantity \[ \eta \] is the fraction of the input that is converted into output.

OpenStudy (anonymous):

im confused what is the answer?